Your company is evaluating alternative plans to finance an expansion of its business as detailed below Your company is considering the best way to finance its new operating division which requires finance of $10m. Plan A involves all equity. Two million new shares will be issued at $5 each. Plan B involves the use of financial leverage. Five million dollars will be raised by selling bonds with a coupon interest rate of 10% p.a. Under this plan, the remaining $5m would be raised from issuing shares at $5 each. The use of financial leverage is considered to be a permanent part of the firm’s capitalisation so no maturity date is needed for the analysis. The company pays tax at the rate of 36% and currently has $1m debt on issue with a coupon interest rate of 10% p.a., and 2m shares. The current financial statements of the firm show earnings before interest and tax (EBIT) of $1,500,000 and earnings per share (EPS) of 41.6 cents. The new project is expected to add $1,200,000 to earnings before interest and tax.

a) Calculate the EBIT indifference point

b) For EBIT = $2,400,000 calculate the EPS for each plan and identify which plan the company should prefer.
Answer #1

Answer :-(PARTA: Previous EBIT (Earning Before Interest & Tones) = 1,500,000 Loh new phoyu евт преди - 1, уро, ооо - Total EB50% equity and son IPLANB Where he finances Deft. EBIT = 2.700,000 Lesson Interest @loy = (1,00,000) Lest Interest @ 101) = ((A) EBIT Ondyperence point Equation & Ý -I) CI-T)-PD = (x-TJ (IT) where x = Point Indifference I - Inleret under allunotive Iexuting Capital structure By company finanuale In the case we are taking and new copital structure - (X-I) (1-T)-PD (X-IQ) (1PARTB Under Plon A) Less EBIT = Interest 2, You ooo C1,00,000) - 2,300,000 less Tax @367 = (8,28,000) Net profit = 1,472,000